How many atoms are in each of the following samples that contain naturally occurring iodine atoms?

(a)127 amu =    atom(s)

(b) 19,436 amu=    atoms

(c) 7.28 x 10^24 amu=    x10^22atoms
Answer #1

(a)127 amu =    atom(s):

127 amu = 127 g / 127 g / mole = 1.0 mole = 6.023 * 10^23 atoms.

thus

number of atoms = 6.023 * 10^23 atoms

(b) 19,436 amu=    atoms

19436 amu = 19436 g / 127 g / mole = 153.04 mole.

number of atoms = 153.04 * 6.023 * 10^23 = 9.22 * 10^25 atoms (answer)

(c) 7.28 x 10^24 amu =    x10^22 atoms

7.28 x 10^24 amu = 7.28 x 10^24 / 127 = 5.73 * 10^22 mole.

number of atoms = 5.73 * 10^22 * 6.023 * 10^23 = 3.45 * 10^46 atoms.

please check your data again.